# A mathematical theory of the critical point of ferromagnetic Ising systems

### v1

@@ 4 Conclusion and open problems
We propose that the general theory in the above remark is the theory of phase transitions (or, equivalently, for quantum spin systems [DLS78]) of spontaneous symmetry breaking (s.s.b.)) in quantum spin systems and relativistic quantum field theory (rqft) (with quantum continuous systems arising from the nonrelativistic limit of the latter). Although in the case of rqft no non-perturbative, experimentally relevant theory (such as gauge theories in four space-time dimensions) exists yet, the two theories have one very important aspect in common: a notion of local dynamics. In the case of quantum spin systems, disturbances propagate with bounded “group velocity”, due to the now famous Lieb-Robinson (LR) bound ([LR72], see also [NSY19] and references given there), in the case of rqft we have the fundamental property of local commutativity [SW64]. Since the LR bound played a crucial role in providing one explanation of irreversibility in ([Wre20], [Wre23]), we may expect that it may be so in connection with the other macroscopic phenomena related to the present paper, discussed in Subsection 1.4.
@@
The main objects in the proposed general theory are microscopic and macroscopic states of infinite systems (this point relates to the second notion of phase transition discussed in Subsection 1.1, and their mutual relationship. The “discontinuities”, mentioned by Glimm as the “essential difference in the analysis over finite vs. infinite dimensional spaces” ([Gli76], p.674), which manifest themselves as singularities in the free energy in the thermodynamic limit, appear also, in this extended view, as changes in basic structural properties of the states of infinite systems, for instance in the change from pure to mixed in the theory of irreversibility in ([Wre20], [Wre23]), or in the passage from mixed to pure states in the theory of superselection sectors originating from interactions with the environment ([NT99], [NW14], see also [Wre23]). In the present paper, the “discontinuity” also arose from taking the limit (1.4.7) in Definition 1.4. Indeed, starting from a (microscopic) equilibrium (KMS) state $\omega_{T}$, the thereby defined (macroscopic) fluctuation state $\omega_{T}^{F}$ may also exhibit basic structural changes with regard to $\omega_{T}$.
In this paper we analysed the Ising universality class (1.2.1), (1.2.2), (1.2.3), but most results concerned the special Ising case $\delta=0$, the case $\delta\neq 0$ and several related issues, in particular concerning the Lie algebra of fluctuations, are found in Subsection 1.4.
One of the main aspects of the mutual relationship between $\omega_{T}$ and $\omega_{T}^{F}$ concerns the relation between the energy gap, for $T=0$, or the rate of clustering of the ergodic components of $\omega_{T}$, for $0\leq T\leq T_{c}$, and certain properties of $\omega_{T}^{F}$; if there is an energy gap at $T=0$, or the clustering is summable ($l^{1}$), at $0<T<T_{c}$, the fluctuations characterizing $\omega_{T}^{F}$ are normal; if the clustering is not summable ($T=T_{c}$), $\omega_{T}^{F}$ has anomalous fluctuations. This was shown in Theorem 2.5 for $\nu=2$ and for $\nu=3$ under the Standard Assumption (2.3.24). The above connections refer only to the connected $2$-point Wightman functions. In the case of s.s.b. of a continuous symmetry (e.g., $\left|\delta\right|=1$ in (1.2.1)), the absence of a gap has been proved in the case of finite range interactions ( [WFW87], [LPW81]) and absence of $l^{1}$- clustering in [LPW81]. Thus, Criterion A does not apply (Proposition 3.1). For long-range interactions, such as the BCS model, s.s.b. of a continuous symmetry is compatible with an energy gap (see [We71] and references given there). In rqft, the two properties – energy gap and exponential decay in space – are related by the Poincaré group, corresponding to a representation of positive mass $m>0$, with the critical case being associated to zero mass representations ($m=0$): the passage from $m>0$ to $m=0$ defines a special contraction of Lie algebra representations ([CW99], [CW79]), see Subsection 1.4. The inclusion of rqft is supported by the former project of Glimm and Jaffe [GJ81] of considering the “quantum field theory of the Ising model”; it is also in close connection with Hegerfeldt’s approach [Heg75] to “prime fields” on Borchers’ tensor algebra. In addition, the theory of quasi-particles (spin-waves) in Subsection 1.4, for $T=0$, has a natural counterpart for nonzero temperature which is different from the zero temperature case in several essential aspects, and is very close in spirit to rqft [NRT83].
@@
The basic element of our framework, following previous work by one of us (M.R.) in [MR02], was to employ a smooth version of fluctuation operators, better adapted to Fourier transform methods. Indeed, a distinction between the critical state and the other ergodic equilibrium states is naturally formulated for the connected two-point Wightman function, in terms of its Fourier transform (2.3.14) which is stronger than the condition of non-$l^{1}$- clustering, but allows a considerable progress towards a proof of the Standard Assumption (Theorem 2.4). Its analogue for the higher ($r\geq 4$)-connected Wightman functions – the Scaling Assumption C ((2.4.2) with $\rho^{{}^{\prime}}>0$ in (2.4.3)) – yields a proof of the quasi-free nature of the fluctuation state $\omega_{T_{c}}^{F}$. Since the corresponding equilibrium state $\omega_{T_{c}}$ is an interacting state, this is a third of the basic structural changes discussed in the second paragraph of this subsection. A different structural change ( see Remark 2.6) was shown to occur in the passage from finite to infinite volume due to Proposition 1.5, which demonstrates that for any finite volume the fluctuation state is, for any temperature, an interacting state. This proposition is a direct corollary of the important work of De Coninck and Gottal [dCG84], which relies on the property of infinite divisibility of the reciprocal of the characteristic function (for finite volume). A further consequence of their results, Proposition 1.6, was used in the proof of Theorem 2.6.
@@
Several aspects of dynamics were shown to be relevant to this general framework. Firstly, a physically intuitive explanation of the quasi-free property of the fluctuation states for all temperatures (given in Remark 1.5 was suggested by the work of Wehrl [We71] on the spin waves in the BCS model, who showed that the fluctuation states do not satisfy the (dynamical, equilibrium) KMS condition. Since the BCS model is a prototype of long-range interactions, and the quasi-free property of the fluctuation state was also proved for short-range interactions in Theorem 2.6, this “non-equilibrium behavior” of the fluctuations does not, apparently, distinguish between the two main universality classes (long-range and short-range). It is therefore natural to investigate what happens with the other macroscopic observables – the intensive observables – in non-equilibrium situations, and whether there a distinction between the two classes may arise. We were able to investigate this question in section 3.2.2. The two universality classes -long-range and short-range – are known to exhibit different rates of decay of unstable states (see Remark 3.2). But, in addition, a more significant aspect was shown for the first time: in the long-range case, different rates of fall-off of the potential yield different rates of decay, i.e., “dynamical critical indices” violate universality. These results, proved in the Dyson model in Subsection 3.2.2 are corollaries of the recent theorem of Albert and Kiessling [AlKi17] on the Cloitre function.
@@
The Bleher-Sinai theory, via hierarchical model, is relevant to the description of the universality classes in the Ising model. It correctly describes the Ising model with long-range interaction and the upper critical dimension of the Ising model [ADC21] and is, in fact, the only way to understand the remarkable results on universality reviewed in Subsection 3.2.1. Their nonlinear equation for the block spin distribution does have a (non-unique, not normal) Gaussian solution, shown to be correct for the Ising model with short-range interactions by Theorem 2.6. It thereby correctly accounts for the anomalous fluctuations, the major result of physical significance. It may be therefore hoped that extensions of the Bleher-Sinai theory to hierarchical models with continuous symmetries (see [MCG08] and references given there) may bring further insights on the basic open problem of what should replace Criterion A in the case of models with s.s.b. of continuous symmetries. In spite of the fact that most magnetic materials are anisotropic, the universality class of the isotropic model (e.g., $\delta=\pm 1$ in (1.2.1)) is certainly not empty![Pel13]. This is one of the major open problems, whose satisfactory solution would be a further hint that the general theory conjectured in this section really exists.
Appendix A Appendix
@@
The result in this appendix provides additional support to the expression “Ising universality class” used in connection with Theorem 1.2, showing that Ising domination occurs in connection with the property of existence of an energy gap. It is contained in [vHBW84], but this reference contains several (sign) misprints which may obscure understanding, and omits certain details. We therefore include a complete proof.
@@
Appendix B Appendix
A simple general probabilistic approach has been proposed by Newman [Ne79]. The main and important parts of that, which are pertinent to the present paper, are now reviewed. Let
@@ Appendix B Appendix
Typically $\sigma_{\Lambda_{n}}=\sum_{x\in\Lambda_{n}}\sigma_{x}$ represents some extensive quantity such magnetization, energy or number of particles contained in a finite box $\Lambda_{n}\subset\mathbb{Z}^{v}$ of volume $\left|\Lambda_{n}\right|=n^{\nu}$ and the expectation $\mathbb{E=E}_{\beta,J,\Lambda_{n}}$ with respect to (1.1.1) omits for simplicity the dependences on its parameters and domains. Associated with $\sigma_{\Lambda_{n}}$, $z$ is the intensive thermodynamic variable such as external magnetic field, inverse temperature or the negative of the chemical potential. Free or periodic conditions are imposed in the Hamiltonian of (1.1.1) at the boundary $\delta\Lambda_{n}$ (with magnetic field omitted and possibly long range interactions $J_{xy}$, $x,y\in\Lambda_{n}$).
@@
Remark B.2.
Concerning hypotheses A. and B., it is of special interest to compare them with what would be the analogous hypotheses in the present paper. This has to do with the choice (B.0.2), which is different in our case. The analogous condition to Hypothesis B. is (1.4.9) with $\omega_{T}$ replaced by $\omega_{\Lambda,T}$, and the characteristic function and its thermodynamic limit given by (1.1.25), (1.1.27), playing the role of $f_{n}(z)$ in (B.0.2), and therefore $f(z)$ in Hypothesis A.; the special case $\Phi_{T}^{\prime}(0)$ in (1.4.9) is the analogue of Hypothesis B., but (1.4.9) requires, in addition, that $\Phi_{T}^{(n)}(0)$ exists for $n=1,2,\ldots$ the superscript $(n)$ denoting derivative of order $n$. This is much stronger than Hypothesis B. because it involves derivatives of all orders, but, much more importantly, (1.1.25) involves the fluctuation operator (1.1.19), with the variance which is anomalous, i.e., satisfies (1.1.21).
The apparent disagreement (Gaussian vs. Non-Gaussian) between our result, Theorem 2.6 and Newman’s (Corollary B.4 of Appendix B), is explained by the fact that our fluctuation state $\omega_{T}^{F}$ is singular with respect to the equilibrium state (see 1.4 and the remarks thereafter), this “singularity” being precisely expressed by the basic structural difference between these states (quasi-free (Gaussian) in the case of $\omega_{T}^{F}$ by Theorem 2.6, and interacting (Non-Gaussian) by Newman’s result (see Corollary B.4 of Appendix B. See also the last paragraph of Remark 1.4 which connects this singularity property with one of the possibilities pointed out by Griffiths and Pearce [GP79] to explain the peculiarities of RG transformation (cf. also van Enter’s review [vE99]). Finally, Remark 1.4 explains why Newman’s results, somewhat surprisingly, actually complement ours in an interesting way.
@@
[…]
@@
According to Lee-Yang theory as it is modernly stated in [Ne75], a random variable $X$ is said of Newman’s type $\mathcal{L}$, if i. there exist some constants $c,C>0$ such that $\left|\mathbb{E}\exp(zX)\right|\leq C\exp\left(c\left|z\right|^{2}\right)$ holds for all $z\in\mathbb{C}$; and ii. the moment generating function $\mathbb{E}\exp(zX)$ is an even function with only pure imaginary zeros. When $X=\sum_{i\in\Lambda_{n}}\sigma_{i}=\sigma_{\Lambda_{n}}$, the so called Lee-Yang property, satisfied for Dyson hierarchical Ising model [HHW01], hierarchical $O(N)$ sigma model [Ko88, Zu94, Wa04] and hierarchical spherical model [BAHM02, MCG08], in combination with the renormalization group (RG) machinery, has allowed to deal with both the strong and weak coupling regimes [HHW01, Wa04, MCG08], starting the RG dynamics from their respective original “a priori” measures, evolving towards the fixed point at the critical point. Although the original goal of Lee-Yang program was to understand phase transitions directly from the distribution of Lee-Yang zeroes, rigorous results are very few ([CJN23, CJN21] and references therein) and restrict to specific models ([CJN23, MCG08] and references therein).

### v2

@@ 4 Conclusion and open problems
In this paper, however, we analysed some aspects of just one universality class, the Ising universality class defined by the equations (1.2.1), (1.2.4), and (1.2.5), and most results concerned only the special Ising case $\delta=0$. This class also includes Euclidean quantum field theory [GJ81].
@@
The main objects in a future general theory might be microscopic and macroscopic states of infinite systems (this point relates to the second notion of phase transition discussed in Subsection 1.1), and their mutual relationship. The “discontinuities”, mentioned by Glimm as the “essential difference in the analysis over finite vs. infinite dimensional spaces” ([Gli76], p.674), which manifest themselves as singularities in the free energy in the thermodynamic limit, might appear also, in this extended view, as changes in basic structural properties of the states of infinite systems, for instance in the change from pure to mixed in the theory of irreversibility in ([Wre20], [Wre23]), or in the passage from mixed to pure states in the theory of superselection sectors originating from interactions with the environment ([NT99], [NW14], see also [Wre23]). In the present paper, the “discontinuity” also arose from taking the limit (1.4.7) in Definition 1.4. Indeed, starting from a (microscopic) equilibrium (KMS) state $\omega_{T}$, the thereby defined (macroscopic) fluctuation state $\omega_{T}^{F}$ may also exhibit basic structural changes with regard to $\omega_{T}$.
One of the main aspects of the mutual relationship between $\omega_{T}$ and $\omega_{T}^{F}$ concerns the relation between the energy gap, for $T=0$, or the rate of clustering of the ergodic components of $\omega_{T}$, for $0\leq T\leq T_{c}$, and certain properties of $\omega_{T}^{F}$; if there is an energy gap at $T=0$, or the clustering is summable ($l^{1}$), at $0<T<T_{c}$, the fluctuations characterizing $\omega_{T}^{F}$ are normal; if the clustering is not summable ($T=T_{c}$), $\omega_{T}^{F}$ has anomalous fluctuations. This was shown in Theorem 2.5 for $\nu=2$ and for $\nu=3$ under the Standard Assumption (2.3.24). The above connections refer only to the connected $2$-point Wightman functions.
In the case of s.s.b. of a continuous symmetry (e.g., $\left|\delta\right|=1$ in (1.2.1)), the absence of a gap has been proved in the case of finite range interactions ( [WFW87], [LPW81]) and absence of $l^{1}$- clustering in [LPW81]. Thus, Criterion A does not apply. For long-range interactions, such as the BCS model, s.s.b. of a continuous symmetry is compatible with an energy gap (see [We71] and references given there).
@@
The basic element of our framework, following previous work by one of us (M.R.) in [MR02], was to emply a smooth version of fluctuation operators, better adapted to Fourier transform methods. Indeed, a distinction between the critical state and the other ergodic equilibrium states is naturally formulated for the connected two-point Wightman function, in terms of its Fourier transform (2.3.14) which is stronger than the condition of non-$l^{1}$- clustering, but allows a considerable progress towards a proof of the Standard Assumption (Theorem 2.4). Its analogue for the higher ($r\geq 4$)-connected Wightman functions – the Scaling Assumption C ((2.4.2) with $\rho^{{}^{\prime}}>0$ in (2.4.3)) – yields a proof of the quasi-free nature of the fluctuation state $\omega_{T_{c}}^{F}$. Since the corresponding equilibrium state $\omega_{T_{c}}$ is an interacting state, this is a third of the basic structural changes discussed in the previous paragraph. A different structural change (see Remark 2.6) was shown to occur in the passage from finite to infinite volume due to Proposition 1.5, which demonstrates that for any finite volume the fluctuation state is, for any temperature, an interacting state. This proposition is a direct corollary of the important work of De Coninck and Gottal [dCG84], which relies on the property of infinite divisibility of the reciprocal of the characteristic function (for finite volume). A further consequence of their results, Proposition 1.6, was used in the proof of Theorem 2.6.
@@
Several aspects of dynamics were shown to be relevant to this general framework. Firstly, a physically intuitive explanation of the quasi-free property of the fluctuation states for all temperatures (given in Remark 1.7 was suggested by the work of Wehrl [We71] on the spin waves in the BCS model, who showed that the fluctuation states do not satisfy the (dynamical, equilibrium) KMS condition. Since the BCS model is a prototype of long-range interactions, and the quasi-free property of the fluctuation state was also proved for short-range interactions in Theorem 2.6, this “non-equilibrium behavior” of the fluctuations does not, apparently, distinguish between the two main universality classes (long-range and short-range). It is therefore natural to investigate what happens with the other macroscopic observables – the intensive observables – in non-equilibrium situations, and whether there a distinction between the two classes may arise. We were able to investigate this question in Subsection 3.2. The two universality classes -long-range and short-range – are known to exhibit different rates of decay of unstable states (see Remark 3.2). But, in addition, a more significant aspect was shown for the first time: in the long-range case, different rates of fall-off of the potential yield different rates of decay, i.e., “dynamical critical indices” violate universality. These results, proved in the Dyson model in Subsection 3.2 are corollaries of the recent theorem of Albert and Kiessling [AlKi17] on the Cloitre function.
@@
The Bleher-Sinai theory, via hierarchical model, is relevant to the description of the universality classes in the Ising model. It correctly describes the Ising model with long-range interaction and the upper critical dimension of the Ising model [ADC21] and is, in fact, the only way to understand the remarkable results on universality reviewed in Subsection 3.2. Their nonlinear equation for the block spin distribution does have a (non-unique, not normal) Gaussian solution, shown to be correct for the Ising model with short-range interactions by Theorem 2.6. It thereby correctly accounts for the anomalous fluctuations, the major result of physical significance.
The following remarks concern the breakdown of universality in certain models which are not in the Ising universality class, but are nevertheless very illuminating regarding the great complexity of the universality issue.
The result of [LPW81] was generalized by Martin [Mar82] in an optimal way to include several types of decay of the interaction potential, in particular the case of dimension $\nu=2$, providing an extension of the Hohenberg-Mermin-Wagner theorem ([H67], [MeW66]). The latter theorem precludes a first-order phase transition (i.e., with nonzero spontaneous magnetization) in n-vector models, with $SO(n)$ symmetry. van Enter and Shlosman [vES02] showed, however, that, in a model with sufficiently nonlinear interaction potentials which satisfies the Hohenberg-Mermin-Wagner theorem, and therefore no spontaneous magnetization takes place, there is a first-order transition between two non-magnetized states. In three or more dimensions the result is that the ordinary XY model has a critical point separating the non-magnetized high temperature regime from the magnetized low temperature regime: in the non-linear case this transition is first-order.
Another non-universal example is the 3-state Potts model. in dimension two it has an ordinary critical point. If the range is large but finite, the transition becomes first-order. This means that the same potts permutation symmetry is broken, both models are finite-range but the behavior is again non-universal ([GM07]).
Finally, it may be therefore hoped that extensions of the Bleher-Sinai theory to hierarchical models with continuous symmetries (see [MCG08] and references given there) may bring further insights on the basic open problem of what should replace Criterion A in the case of models with s.s.b. of continuous symmetries. In spite of the fact that most magnetic materials are anisotropic, the universality class of the isotropic model (e.g., $\delta=\pm 1$ in (1.2.1)) is certainly not empty![Pel13]. This is one of the major open problems, whose satisfatory solution would be a further hint that the general theory conjectured in this section really exists.
Appendix A “Ising universality class”
@@
The result in this appendix provides additional support to the expression i the title, which is used in connection with Theorem 1.2, showing that Ising domination occurs in connection with the property of existence of an energy gap. It is contained in [vHBW84], but this reference contains several (sign) misprints which may obscure understanding, and omits certain details. We therefore include a complete proof.
@@
Appendix B Possible non Gaussian behavior at $\nu=4$
A simple general probabilistic approach has been proposed by Newman [Ne79], of which we present a short review.
It is complementary to the present paper because it concerns the dimension $\nu=4$ in the critical case, or $\nu=3$ in the tricritical case.
Let
@@ Appendix B Possible non Gaussian behavior at $\nu=4$
Tipically $\sigma_{\Lambda_{n}}=\sum_{x\in\Lambda_{n}}\sigma_{x}$ represents some extensive quantity such magnetization, energy or number of particles contained in a finite box $\Lambda_{n}\subset\mathbb{Z}^{v}$ of volume $\left|\Lambda_{n}\right|=n^{\nu}$ and the expectation $\mathbb{E=E}_{\beta,J,\Lambda_{n}}$ with respect to (1.1.1) omits for simplicity the dependences on its parameters and domains. Associated with $\sigma_{\Lambda_{n}}$, $z$ is the intensive thermodynamic variable such as external magnetic field, inverse temperature or the negative of the chemical potential. Free or periodic conditions are imposed in the Hamiltonian of (1.1.1) at the boundary $\delta\Lambda_{n}$ (with magnetic field omitted and possibly long range interactions $J_{xy}$, $x,y\in\Lambda_{n}$).
@@
$f_{e}(z)$$\leq$ $K\left|z\right|^{p}\left|\log\left|z\right|\right|^{\theta}\ ,\qquad\text{for }z\ \text{small}$
$\tau_{n}^{2}$$\geq$ $\varepsilon n^{\nu+2-\eta}\left|\log n\right|^{\psi}\ ,\qquad\text{for }n\ \text{large}$ (B.0.15)
@@
According to Lee-Yang theory as it is modernly stated in [Ne75], a random variable $X$ is said of Newman’s type $\mathcal{L}$, if i. there exist some constants $c,C>0$ such that $\left|\mathbb{E}\exp(zX)\right|\leq C\exp\left(c\left|z\right|^{2}\right)$ holds for all $z\in\mathbb{C}$; and ii. the moment generating function $\mathbb{E}\exp(zX)$ is an even function with only pure imaginary zeros. When $X=\sum_{i\in\Lambda_{n}}\sigma_{i}=\sigma_{\Lambda_{n}}$, the so called Lee-Yang property, satisfied for Dyson hierarchical Ising model [HHW01], see also [CE78], hierarchical $O(N)$ sigma model [Ko88, Zu94, Wa04] and hierarchical spherical model [BAHM02, MCG08], in combination with the renormalization group (RG) machinery, has allowed to deal with both the strong and weak coupling regimes [HHW01, Wa04, MCG08], starting the RG dynamics from their respective original “a priori” measures, evolving towards the fixed point at the critical point. Although the original goal of Lee-Yang program was to understand phase transitions directly from the distribution of Lee-Yang zeroes, rigorous results are very few ([CJN23, CJN21] and references therein) and restrict to specific models ([CJN23, MCG08] and references therein).
@@ Acknowledgements
We are, above all, grateful to the referee, who advised us to eliminate a large part of irrelevant material, and concentrate only in the Ising universality class. This led us to a streamlined version, in which the open problems are clearly separated from the true subject of the paper. We also thank him for several important remarks, which contributed to render the discussion completely precise, and considerably enriched the paper with observations, mostly regarding the concept of universality.